OpenStudy (aimmebgood):

-2y^(2) x -5y would it come out as 10y^(2) or ^(3) or something else?

Parth (parthkohli):

Yeah, ^(3)

hartnn (hartnn):

-2*-5 = 10 and y^2*y = y^3 so 10y^3

Parth (parthkohli):

\[-2y^2 \times -5y\]\[\implies -2 \times y \times y \times -5 \times y\]\[\implies (-2 \times -5) \times (y \times y \times y)\]